Find the factored form of y = 2x^2 + 6x + 4
2(x - 1)(x - 2)
(x + 1)(x + 2)
2(x - (-1))(x - (-2))
(x - (-1))(x - (-2))
Find the factored form of y = x^2 + 4
(2x - i)(2x + i)
(x - 2)(x - (-2))
(x - 2i)(x - (-2i))
(x - 0)(x - 4)
Find the usual y = ax^2 + bx + c form of the quadratic which has a = 1, and two zeros z = 3 and z' = 7
y = x^2 - 10x + 21
y = x^2 + 10x - 21
y = x^2 + 3x - 7
y = x^2 - 3x + 7
Find the usual y = ax^2 + bx + c form of the quadratic which has a = 1, and two zeros z = 2 + i and z' = 2 - i
y = x^2 - 4x + 5
y = x^2 + 4x + 5
y = x^2 + 4x - 5
y = x^2 + 4






Answer by stanbon(75887)

Find the factored form of y = 2x^2 + 6x + 4
Answer:
(x - (-1))(x - (-2))
Find the factored form of y = x^2 + 4
Answer:
(x - 2i)(x - (-2i))
Find the usual y = ax^2 + bx + c form of the quadratic which has a = 1, and two zeros z = 3 and z' = 7
Answer:
y = x^2 - 10x + 21
Find the usual y = ax^2 + bx + c form of the quadratic which has a = 1, and two zeros z = 2 + i and z' = 2 - i
Answer:
y = x^2 - 4x + 5
--------------
Cheers,
Stan H.